# A universal speed limit for spreading of coherence

### v3

@@ Abstract
Discoveries of fundamental limits for the rates of physical processes, from the speed of light to the Lieb–Robinson bound for information propagation [1, 2, 3, 4], are conceptual breakthroughs that often challenge our understanding of the underlying physics. Here we observe such a limit for a paradigmatic many-body phenomenon, the spreading of coherence during formation of a weakly interacting Bose–Einstein condensate [5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21]. We study condensate formation in an isolated homogeneous atomic gas [22, 23] that is initially far from equilibrium, in an incoherent low-energy state, and condenses as it relaxes towards equilibrium. Tuning the inter-atomic interactions that drive condensation, we show that the spreading of coherence through the system is initially slower for weaker interactions, and faster for stronger ones, but always eventually reaches the same limit, where the square of the coherence length grows at a universal rate given by the ratio of Planck’s constant and the particle mass, or equivalently by the quantum of velocity circulation associated with a quantum vortex. These observations are robust to changes in the initial state, the gas density, and the system size. Our results provide benchmarks for theories of universality far from equilibrium [24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36], are relevant for quantum technologies that rely on large-scale coherence, and invite similar measurements in other systems.
@@
Understanding the dynamics of far-from-equilibrium many-body systems, including the emergence of long-range order in such systems, is an outstanding problem in physics, relevant from subnuclear to cosmological lengthscales [24, 25, 27, 26, 28, 29, 30, 31, 32, 33, 34, 35, 36]. Conceptually, far-from-equilibrium relaxation and emergence of coherence have long been linked to decaying turbulence [25, 14, 28], which often features self-similar scaling dynamics. More recently, within the framework of nonthermal fixed points (NTFPs) [29], theorists have drawn analogies between such dynamics and the equilibrium properties of systems close to a continuous phase transition. Near the transition to an ordered state of matter, such as a superfluid or a ferromagnet, the system is scale invariant and its salient properties do not depend on the microscopic details [37]. Analogously, in the NTFP theory, far-from-equilibrium systems, including the early universe undergoing reheating [28], quark-gluon plasma in heavy-ion collisions [38], quantum magnets [39], and ultracold atomic gases [40, 41, 42, 43, 44, 45], generically display dynamic (spatiotemporal) scaling, with scaling exponents that could define far-from-equilibrium universality classes. Recently, far-from-equilibrium dynamic scaling was observed in several experiments with ultracold atoms, in both isolated (relaxing) [46, 47, 48, 49, 50, 51, 52, 53] and continuously driven [54, 55] systems.
@@
Here, we go beyond the elegant scaling properties of far-from-equilibrium relaxation [46, 47, 48, 49, 50, 51, 52, 53], to address the crucial question of how long it takes to establish long-range order. We study this problem for the paradigmatic macroscopically coherent state, the weakly interacting Bose–Einstein condensate, which is also a textbook example of a superfluid.
@@
with $\beta=1/2$ [41, 43, 36]. In our weakly interacting gas, with the vortex-core size set by the healing length $\xi=1/\sqrt{8\pi na}$, both vortices and waves could play a significant role. Note, however, that the predictions of Eqs. (1) and (2) differ only in a logarithmic correction. Moreover, our measurements of $n_{k}$ probe $\ell$ independently of what type of excitations are dominant and limit its value.
@@
Specifically for $\beta=1/2$, as observed in Fig. 1, this has a counter-intuitive implication that $D=\textrm{d}\ell^{2}/\textrm{d}t\sim\hbar/m$ does not depend on the interaction strength $na$ (see also [60]). In the hydrodynamic theory of Ref. [25], the same result emerges if one neglects the logarithmic correction in Eq. (1) (which depends on the interactions through $\xi$), because then $\textrm{d}\ell^{2}/\textrm{d}t$ can only be set by the quantum of velocity circulation associated with a quantum vortex, $\kappa=2\pi\hbar/m$.
@@
We thank Andrey Karailiev, Maciej Gałka, Timon Hilker, Mikhail Lukin, Robert Smith, and Tobias Donner for useful discussions and comments on the manuscript. This work was supported by EPSRC [Grant No. EP/P009565/1], ERC (UniFlat), and STFC [Grants No. ST/T006056/1 and No. ST/Y004469/1]. C. E. acknowledges support from Jesus College (Cambridge). Z. H. acknowledges support from the Royal Society Wolfson Fellowship.
Data availability The data that support the findings of this study are available in the Apollo repository [url to be added]. Any additional information is available from the corresponding authors upon a reasonable request.

### v4

@@ Abstract
Discoveries of fundamental limits for the rates of physical processes, from the speed of light to the Lieb–Robinson bound for information propagation [1, 2], often lead to breakthroughs in the understanding of the underlying physics. Here we observe such a limit for a paradigmatic many-body phenomenon, the spreading of coherence during formation of a weakly interacting Bose–Einstein condensate [3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18]. We study condensate formation in an isolated homogeneous atomic gas [19, 20] that is initially far from equilibrium, in an incoherent low-energy state, and condenses as it relaxes towards equilibrium. Tuning the inter-atomic interactions that drive condensation, we show that the spreading of coherence through the system is initially slower for weaker interactions, and faster for stronger ones, but always eventually reaches the same limit, where the square of the coherence length grows at a universal rate given by the ratio of Planck’s constant and the particle mass, or equivalently by the quantum of velocity circulation associated with a quantum vortex. These observations are robust to changes in the initial state, the gas density, and the system size. Our results provide benchmarks for theories of universality far from equilibrium [21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34], are relevant for quantum technologies that rely on large-scale coherence, and invite similar measurements in other systems.
@@
Understanding the dynamics of far-from-equilibrium many-body systems, including the emergence of long-range order in such systems, is an outstanding problem in physics, relevant from subnuclear to cosmological lengthscales [21, 22, 23, 24, 26, 25, 27, 28, 29, 30, 31, 32, 33, 34]. Conceptually, far-from-equilibrium relaxation and emergence of coherence have long been linked to decaying turbulence [21, 23, 12, 27], which often features self-similar scaling dynamics. More recently, within the framework of nonthermal fixed points (NTFPs) [28], theorists have drawn analogies between such dynamics and the equilibrium properties of systems close to a continuous phase transition. Near the transition to an ordered state of matter, such as a superfluid or a ferromagnet, the system is scale invariant and its salient properties do not depend on the microscopic details [35]. Analogously, in the NTFP theory, far-from-equilibrium systems, including the early universe undergoing reheating [27], quark-gluon plasma in heavy-ion collisions [36], quantum magnets [37], and ultracold atomic gases [38, 39, 40, 41, 42, 43], generically display dynamic (spatiotemporal) scaling, with scaling exponents that could define far-from-equilibrium universality classes. Recently, far-from-equilibrium dynamic scaling was observed in several experiments with ultracold atoms, in both isolated (relaxing) [44, 45, 46, 47, 48, 49, 50] and continuously driven [51, 52] systems.
@@
Here, we go beyond the elegant scaling properties of far-from-equilibrium relaxation [44, 45, 46, 47, 48, 49, 50], to address the crucial question of how long it takes to establish long-range order. We study this problem for the paradigmatic macroscopically coherent state, the weakly interacting Bose–Einstein condensate, which is also a textbook example of a superfluid.
@@
with $\beta=1/2$ [21, 39, 41, 34]. In our weakly interacting gas, with the vortex-core size set by the healing length $\xi=1/\sqrt{8\pi na}$, both vortices and waves could play a significant role. Note, however, that the predictions of Eqs. (1) and (2) differ only in a logarithmic correction. Moreover, our measurements of $n_{k}$ probe $\ell$ independently of what type of excitations are dominant and limit its value.
@@
Specifically for $\beta=1/2$, as observed in Fig. 1, this has a counter-intuitive implication that $D=\textrm{d}\ell^{2}/\textrm{d}t\sim\hbar/m$ does not depend on the interaction strength $na$ (see also [21, 24]). In the hydrodynamic theory of Ref. [23], the same result emerges if one neglects the logarithmic correction in Eq. (1) (which depends on the interactions through $\xi$), because then $\textrm{d}\ell^{2}/\textrm{d}t$ can only be set by the quantum of velocity circulation associated with a quantum vortex, $\kappa=2\pi\hbar/m$.
@@
We thank Andrey Karailiev, Maciej Gałka, Timon Hilker, Mikhail Lukin, Robert Smith, Tobias Donner, Gregory Eyink, and Vladimir Rosenhaus for useful discussions and comments on the manuscript. This work was supported by EPSRC [Grants No. EP/P009565/1 and EP/Y01510X/1], ERC (UniFlat), and STFC [Grants No. ST/T006056/1 and No. ST/Y004469/1]. C. E. acknowledges support from Jesus College (Cambridge). Z. H. acknowledges support from the Royal Society Wolfson Fellowship.